An electrician experienced a mild shock when he accidentally touched a wire carrying 5.0 A for 1.5 s.



a.) How many coulombs of charge flow through the electrician?



b.) How many electrons constitute the given current?

1
Expert's answer
2022-05-28T08:23:36-0400

(a)

"q=It\\=\rm 5.0\: A\times 1.5\: s=7.5\; C"

(b)

"N=\frac{q}{e}\\
=\frac{7.5}{1.6*10^{-19}}=4.7*10^{19}"